# Politecnico di Torino, DIMEAS, Corso Duca degli Abruzzi, 24 10129 Torino, Italy

### v1

@@ IV.2.2 Long-term statistical predictions
Figure 10: Lorenz-63 system, closed-loop statistical predictions. The top panels present the result of the true and classical (CRC) network. The bottom panels compare the true with the QRC predictions. The shaded region in both figures indicates the uncertainty associated with different random seeds in both CRC and QRC.
@@ IV.3 Higher-dimensional Lorenz-96 model
In this section, we extend our comparison of classical and quantum reservoirs to higher-dimensional chaotic systems. We follow the same procedure defined in Sec. IV.2 and analyse the Lorenz-96 model Lorenz (1995).
@@
where $F$ is the external body forcing term which we set to $F$ = 8, to ensure chaotic behavior Vlachas et al. (2020). We apply periodic boundary conditions, i.e. $x_{1}=x_{m+1}$, and study the reduced order model of Lorenz-96 with ten dimensions ($m=10$). The training set covers an evolution time of $200$ LT for the Lorenz-96 system with a Leading Lyapunov Exponent value of $\Lambda_{1}=1.2$. In Tab. 3, we present the hyperparameters for the ten-dimensional Lorenz-96 system. The set of hyperparameters is derived by using the same procedure as discussed in Sec. II.
@@ IV.3.1 Time-accurate predictions
We analyse the time-accurate prediction capabilities of CRC and different QRC architectures for a ten-dimensional reduced-order model of Lorenz-96. Fig. 11 shows the VPT values of both networks. The QRC architectures C1, C2, and C4 outperform the classical reservoir for increasing reservoir sizes. To study the robustness of the model, the results are averaged over five different sets of optimal hyperparameters and twenty different starting points. For a particular realization of hyperparameters and a starting point, at reservoir size 4096 the VPT value for CRC can be as large as $\sim$ 10 LT. However, the performance is not robust and degrades quickly for other hyperparameter choices and different initial points. Improved performance scaling and robustness indicate a potential quantum advantage in quantum reservoir computers.
@@
Figure 12: Lorenz-96 system with ten dimensions. VPT/circuit depth vs. reservoir sizes for different quantum reservoir configurations.
Similar to the results in the Lorenz-63 system, the choice of a feature map and quantum reservoir architectures is critical for the performance. In Fig. 11, QRC-C4 and C1 have comparable VPT values at different reservoir sizes. Whereas QRC-C2 underperforms by a small factor, QRC-C3 and C5 do not improve with the increasing reservoir size due to not having a recurrent feedback loop. In Fig. 12, QRC-C4 achieves the highest VPT to circuit depth ratio for all of the reservoir sizes. This is due to a lack of recurrence. In comparison, QRC-C1 and C2 require exponentially higher circuit depths at increasing reservoir sizes to achieve similar performance.
@@ IV.3.2 Long-term statisitical predictions
In Figs. 13 and 14, we compare the long-term statistical predictions. The displayed results show the best-performing classical reservoir and quantum reservoir network (QRC-C4) outputs. We evolve these networks for 400 LT in a closed-loop, for the reservoir size of 1024. Both quantum and classical reservoir networks can recover long-term statistics of the chaotic model. The shaded region highlights the uncertainty of different realizations. We find that QRC-C4 predicts the statistical variables with smaller uncertainties when compared to the classical reservoir with equal reservoir sizes.
@@
Figure 13: Ten dimensional Lorenz-96 system. Closed-loop statistical predictions using CRC. The shaded region indicates the uncertainty associated with different random seeds in CRC.
@@
Figure 14: Ten dimensional Lorenz-96 system. Closed-loop statistical predictions using QRC. The shaded region indicates the uncertainty associated with different random seeds in QRC.
IV.4 Chaotic Shear Flow Model for Extreme Event Forecasting
@@
In fluid mechanics, extreme events are sudden and unmitigated changes of observables. The forecasting of extreme events is the first step towards the control and suppression of these violent bursts. Recurrent Neural Networks and reservoir computers are used to study these events qualitatively and quantitatively Racca and Magri (2022b); Srinivasan et al. (2019); Doan et al. (2019a). We consider a qualitative low-order model of turbulent shear flows, which is based on Fourier modes and describes a self-sustained turbulent process. This model is also known as the ’Moehlis, Faisst, and Eckhardt (MFE)’ model. The MFE model is nonlinear and it captures the relaminarization and turbulent bursts Moehlis et al. (2004). Owing to the nonlinear nature of this model, the MFE model has been employed to study turbulence transitions and chaos predictability noa (a); Srinivasan et al. (2019). Mathematically, the MFE system can be described by the non-dimensional Navier-Stokes equations for forced incompressible flow
$\frac{d\boldsymbol{v}}{dt}=-(\boldsymbol{v}\,.\,\boldsymbol{\nabla})\,\boldsymbol{v}\,-\boldsymbol{\nabla}\,{p}+\frac{1}{Re}\Delta\boldsymbol{v}+\boldsymbol{F}(y),\quad\quad\quad\boldsymbol{\nabla}.\boldsymbol{v}=0$ (14)
@@
where $\boldsymbol{v}=(u,v,w)$ is the three-dimensional velocity vector, $p$ is the pressure, $Re$ is the Reynolds number, $\nabla$ is the gradient, and $\Delta$ is the Laplacian operator. $\boldsymbol{F}(y)$ on the right-hand side is the sinusoidal body forcing term and equals $\boldsymbol{F}(y)=\sqrt{2}\pi^{2}/(4Re)\,\text{sin}(\pi y/2)\,\boldsymbol{e_{x}}$. The body forcing term is applied along the $x,y$ direction of the shear between the plates. We consider a three-dimensional domain of length $L_{x},\,L_{y},\,L_{z}$ = $[4\pi,2,2\pi]$ and apply free slip boundary conditions at $y=L_{y}/2$, periodic boundary conditions at $x=[0;L_{x}]$ and $z=[0;L_{z}]$. The set of PDEs can be converted into ODEs by projecting the velocities onto Fourier modes as given by Eq. (15)
$\boldsymbol{v}(\boldsymbol{x},t)=\sum_{i=1}^{9}\,a_{i}(t)\,\hat{\boldsymbol{v}_{i}}\,(\boldsymbol{x}).$ (15)
@@
These nine decompositions for the amplitudes ${a_{i}(t)}$ are substituted into Eq. (14) to yield a set of nine ordinary differential equations as in Moehlis et al. (2004). The MFE system displays a chaotic transient, which in the long term converges to a stable laminar solution. We want to predict the turbulent burst of kinetic energy and chaotic transients, which are extreme events. Fig. 15 shows the evolution of kinetic energy ($k(t)=\frac{1}{2}\sum_{i=1}^{9}\,a_{i}^{2}(t)\,$) and the associated extreme event, and the long-term statistical distribution of the kinetic energy. An extreme event occurs when the kinetic energy $k(t)$ value exceeds the threshold ($k(t)\geq k_{e})$. In this work, we take $k_{e}=0.1$, as in Racca and Magri (2022b)
@@
We solve the MFE system ODEs using an RK4 solver with $dt=0.25s$. The leading Lyapunov Exponent is $\Lambda=0.0163$ for MFE model Racca and Magri (2022b). Unlike the Lorenz-63 and Lorenz-96 models, generating a single long-time series for washout, training, and test sets is not feasible for MFE because a single long-time series eventually laminarizes and gives limited information about the chaotic transients. To address this issue, first, we generate an ensemble of 2000 time series from different random initial points. We take the length of a single time series equal to 65 LT or 16000 time-steps. Second, we discard the time series whose maximum kinetic energy is larger than the laminarization threshold $k_{l}=0.48$. This threshold is selected to be close to the (asymptotic) laminarization value of, $k=0.5$. Out of the 2000 time series, $27\%$ of the series laminarizes. We divide the remaining 1441 time series into washout, training, and test sets. We train our network on a 25-time series, each of length $20LT$. The test set consists of 500 time series. In Sections IV.4.1 and IV.4.2, we compare the time-accurate and statistical predictions for the classical reservoir method against the best-performing quantum reservoir QRC-C4.
@@ IV.4.1 Time-accurate predictions
We use a metric to quantify extreme event prediction capabilities Predictability Horizon (PH) Racca and Magri (2022b); noa (a); Pathak et al. (2018), which is defined as the time interval during which the predicted kinetic energy $k_{pred}$ and true kinetic energy $k_{true}$ are bounded by
@@
where $k_{e}$ is the extreme event threshold, which is equal to 0.1 here as in Racca and Magri (2022b). $\bar{k}$ is the time average of the kinetic energy, and 0.2 is the user-defined error threshold Racca and Magri (2022b). To quantify PH, we sample 100 different extreme events from the unseen test set data. For each extreme event, we start the closed-loop predictions 12 LT before the extreme event and discard the transients in the washout phase for 2 LT. This gives the $\delta_{k_{e}}$ equal to 10 LT. The value of $\delta_{k_{e}}$ represents the time difference from the start of closed-loop predictions to the extreme event. If the PH value is greater than $\delta_{k_{e}}$ we take PH equals to $\delta_{k_{e}}$. If the prediction diverges before the extreme event occurs, or the PH value is less than $\delta_{k_{e}}$, we decrease the $\delta_{k_{e}}$ value by a factor of $\tau_{e}$. We take $\tau_{e}$ = 0.5 LT and the new $\delta_{k_{e}}$ value equals to 9.5 LT. For the next step, we repeat the closed-loop predictions for the extreme event occurring at 9.5 LT i.e. $\delta_{k_{e}}$ = 9.5 LT. We repeat the same method decreasing $\delta_{k_{e}}$ by $\tau_{e}$ until the PH is greater than $\delta_{k_{e}}$ (Fig. 16). That is, the extreme event is predicted correctly by the reservoir computers by the value of PH in the future.
@@
We calculate the PH for a set of reservoir sizes for both quantum and classical reservoir networks. For a classical reservoir, the number of neurons is varied from $N=256$ to $N=2048$. We tune the hyperparameter for each reservoir and calculate the corresponding PH value. We also perform the same analysis for the quantum reservoir with qubits $n=8$ to $n=11$, which, in principle, corresponds to the classical reservoirs of $N=256$ to $N=2048$. The performance of the quantum reservoir also improves with the reservoir size. In Fig. 17, the comparison of classical and quantum reservoirs for various reservoir sizes is shown. Both reservoir networks have comparable performances for a reservoir sizes $N=\{256,512\}$, with the classical reservoir out-performing quantum reservoir by a small margin. However, the PH value of the classical reservoir converges with a median value of around 5.5 LT. This convergence of PH value with increasing reservoir sizes is due to performance saturation in CRC and it is consistent with the results in Racca and Magri (2022b). In contrast, the quantum reservoir outperforms the classical reservoir for increasing reservoir sizes and has a median PH value of 8.09LT with 11 qubits.
A second statistical measure to assess the extreme event prediction capabilities is the $F$-Score Goutte and Gaussier (2005), which measures the model’s accuracy by combining metrics such as precision and recall. To calculate the $F$-score, we measure three different occurrences: (a) an extreme event occurs in the true data set and the model also predicts the extreme event (True Positive, TP) (b) there is no extreme event in the true data but the model predicts an extreme event (False Positive, FP) (c) an extreme event occurs in the true data set but the model does not predict an extreme event, (False Negative, FN) Racca and Magri (2022b). These three metrics are combined to calculate Precision ($p$), Recall ($r$), and $F$-score ($F$) from 1000 different starting points, which span 20 different time series
@@
Figure 17: MFE system, Predictability Horizon (PH) for predictions with different reservoir sizes.
Figure 18: MFE system. $75^{th}$ percentile of Precision , Recall, F-score Vs. On the left side, Prediction Time (PT) for a constant reservoir size $(N=1024)$. On the right side, Reservoir Size for a constant Prediction Time $(PT=3LT)$.
Second, we fix the prediction time to 3 LT (PT = 3LT), which indicates the model is predicting extreme events between 3 - 4 LT. After fixing the PT value, we vary the reservoir sizes from $N=256$ to $N=2048$ and compare the performance metrics. The performances are similar for both quantum and classical reservoirs with reservoir sizes of $\{256,512\}$. Similar to the results shown for the predictability horizon previously, the quantum architecture has a higher $F$-score value for 10,11 qubits as compared to the classical reservoirs of the same sizes (Fig. 18). All the results are computed for 5 different ensemble of networks and for 1000 different starting points.

### v2

@@ IV.2.2 Long-term statistical predictions
Figure 11: Lorenz-63 system. Closed-loop statistical predictions. The top panels present the result of the true and classical (CRC) network. The bottom panels compare the true with the QRC predictions. The shaded region in both figures indicates the uncertainty associated with different random seeds in both CRC and QRC.
@@ IV.3 Higher-dimensional Lorenz-96 model
In this section, we extend our comparison of classical and quantum reservoirs to higher-dimensional chaotic systems. We follow the same procedure defined in Sec. IV.2 and analyze the Lorenz-96 model [36].
@@
where $F$ is the external body forcing term which we set to $F$ = 8, to ensure chaotic behavior [9]. We apply periodic boundary conditions, i.e. $x_{1}=x_{m+1}$, and study the reduced order model of Lorenz-96 with ten dimensions ($m=10$). The training set covers an evolution time of $200$ LT for the Lorenz-96 system with a leading Lyapunov exponent value of $\Lambda_{1}=1.2$. In Tab. 3, we present the hyperparameters for the ten-dimensional Lorenz-96 system. The set of hyperparameters is derived by using the same procedure as discussed in Sec. II.
@@ IV.3.1 Time-accurate predictions
We analyze the time-accurate prediction capabilities of CRC and different QRC architectures for a ten-dimensional reduced-order model of Lorenz-96. Fig. 12 shows the VPT values of both networks. The QRC architectures C1, C2, and C4 outperform the classical reservoir for increasing reservoir sizes. To study the robustness of the model, the results are averaged over five different sets of optimal hyperparameters and 20 different starting points. For a particular realization of hyperparameters and a starting point, at reservoir size 4096 the VPT value for CRC can be as large as $\sim$ 10 LT. However, the performance is not robust and degrades quickly for other hyperparameter choices and different initial points. Improved performance scaling and robustness indicate a potential quantum advantage in quantum reservoir computers.
@@
Similar to the results in the Lorenz-63 system, the choice of a feature map and quantum reservoir architectures is critical for the performance. In Fig. 12, QRC-C4 and C1 have comparable VPT values at different reservoir sizes while QRC-C2 underperforms slightly. The performance of QRC-C3 and C5 do not improve with the reservoir size because there is no feedback loop, and because the linear input feature map only utilises a fraction of state space. By contrast, QRC-C4 (RF-QRC), which also does not have a recurrent feedback loop, has a fully connected input feature map that enriches the reservoir dynamics and has a comparable VPT to QRC-C1 and C2.
As compared to QRC with recurrent connections, QRC-C4 achieves the highest VPT to circuit depth ratio for all of the reservoir sizes as shown in Fig. 13. This is due to a lack of recurrence. In comparison, QRC-C1 and C2 require exponentially higher circuit depths at increasing reservoir sizes to achieve similar performance.
Figure 13: Lorenz-96 system with ten dimensions. VPT/circuit depth vs reservoir sizes for different quantum reservoir configurations.
@@ IV.3.2 Long-term statisitical predictions
In Figs. 14 and 15, we compare the long-term statistical predictions. The displayed results show the best-performing classical reservoir and quantum reservoir network (QRC-C4) outputs. We evolve these networks for 400 LT in a closed-loop, for the reservoir size of 1024. Both quantum and classical reservoir networks can recover long-term statistics of the chaotic model. The shaded region highlights the uncertainty of different realizations. We find that QRC-C4 predicts the statistical variables with smaller uncertainties when compared to the classical reservoir with equal reservoir sizes.
@@
Figure 14: Lorenz-96 system with ten dimensions. Closed-loop statistical predictions using CRC. The shaded region indicates the uncertainty associated with different random seeds in CRC.
@@
Figure 15: Lorenz-96 system with ten dimensions. Closed-loop statistical predictions using QRC. The shaded region indicates the uncertainty associated with different random seeds in QRC.
IV.4 Chaotic shear flow model for extreme events forecasting
@@
In fluid mechanics, extreme events are sudden and unmitigated changes of observables. The forecasting of extreme events is the first step towards the control and suppression of these violent bursts. Recurrent Neural Networks and reservoir computers are used to study these events qualitatively and quantitatively [18, 7, 8]. We consider a qualitative low-order model of turbulent shear flows, which is based on Fourier modes and describes a self-sustained turbulent process. This model is also known as the MFE model. The MFE model is nonlinear and it captures the relaminarization and turbulent bursts [37]. Owing to the nonlinear nature of this model, the MFE model has been employed to study turbulence transitions and chaos predictability [19, 7]. Mathematically, the MFE system can be described by the non-dimensional Navier-Stokes equations for forced incompressible flow
$\frac{{\partial}\pmb{v}}{{\partial}t}=-(\pmb{v}\,.\,\pmb{\nabla})\,\pmb{v}\,-\pmb{\nabla}\,{p}+\frac{1}{Re}\Delta\pmb{v}+\pmb{F}(y),\quad\quad\pmb{\nabla}.\pmb{v}=0$ (14)
@@
where $\pmb{v}=(u,v,w)$ is the three-dimensional velocity vector, $p$ is the pressure, $Re$ is the Reynolds number, $\nabla$ is the gradient, and $\Delta$ is the Laplacian operator. $\pmb{F}(y)$ on the right-hand side is the sinusoidal body forcing term and equals $\pmb{F}(y)=\sqrt{2}\pi^{2}/(4Re)\,\text{sin}(\pi y/2)\,\pmb{e_{x}}$. The body forcing term is applied along the $x,y$ direction of the shear between the plates. We consider a three-dimensional domain of length $L_{x},\,L_{y},\,L_{z}$ = $[4\pi,2,2\pi]$ and apply free slip boundary conditions at $y=L_{y}/2$, periodic boundary conditions at $x=[0;L_{x}]$ and $z=[0;L_{z}]$. The set of PDEs can be converted into ODEs by projecting the velocities onto Fourier modes as given by Eq. (15)
$\pmb{v}(\pmb{x},t)=\sum_{i=1}^{9}\,a_{i}(t)\,\hat{\pmb{v}_{i}}\,(\pmb{x}).$ (15)
@@
These nine decompositions for the amplitudes ${a_{i}(t)}$ are substituted into Eq. (14) to yield a set of nine ordinary differential equations as in Ref. [37]. The MFE system displays a chaotic transient, which in the long term converges to a stable laminar solution. We want to predict the turbulent burst of kinetic energy and chaotic transients, which are extreme events. Fig. 16 shows the evolution of kinetic energy [$k(t)=\frac{1}{2}\sum_{i=1}^{9}\,a_{i}^{2}(t)\,$] and the associated extreme event, and the long-term statistical distribution of the kinetic energy. An extreme event occurs when the kinetic energy $k(t)$ value exceeds the threshold ($k(t)\geq k_{e})$. In this work, we take $k_{e}=0.1$, as in Ref. [18].
@@
We solve the MFE system ODEs using an RK4 solver with $dt=0.25s$. The leading Lyapunov exponent is $\Lambda=0.0163$ for MFE model [18]. Unlike the Lorenz-63 and Lorenz-96 models, generating a single long-time series for washout, training, and test sets is not feasible for MFE because a single long-time series eventually laminarizes and gives limited information about the chaotic transients. To address this issue, first, we generate an ensemble of 2000 time series from different random initial points. We take the length of a single time series equal to 65 LT or 16000 time steps. Second, we discard the time series whose maximum kinetic energy is larger than the laminarization threshold $k_{l}=0.48$. This threshold is selected to be close to the (asymptotic) laminarization value of $k=0.5$. Out of the 2000 time series, $27\%$ of the series laminarizes. We divide the remaining 1441 time series into washout, training, and test sets. We train our network on a 25-time series, each of length $20LT$. The test set consists of 500 time series. In Secs.  IV.4.1 and IV.4.2, we compare the time-accurate and statistical predictions for the classical reservoir method against the best-performing quantum reservoir QRC-C4.
Table 4: Parameters for the MFE system.
Parameters Symbol Classical RC Quantum RC
Time step $dt$ 0.25s 0.25s
Input scaling $\sigma_{in}$ [0 , 1] -
Spectral radius $\rho$ [0.1 , 1] 1
Leak rate $\epsilon$ [0.05, 1] [0.05, 0.3]
Resevoir density $D$ 0.1,0.6,0.9 Configuration C4 (RF-QRC)
@@ IV.4.1 Time-accurate predictions
To quantify extreme event prediction capabilities, we use the Predictability Horizon (PH) [18, 19, 56], which is defined as the time interval during which the predicted kinetic energy $k_{pred}$ and true kinetic energy $k_{true}$ are bounded by
@@
where $k_{e}$ is the extreme event threshold, which is equal to 0.1 here as in Ref. [18]. $\bar{k}$ is the time average of the kinetic energy, and 0.2 is the user-defined error threshold [18]. To quantify the PH, we sample 100 different extreme events from the unseen test set data. For each extreme event, we start the predictions 12 LT before the extreme event and discard the transients in the washout phase for 2 LT by running reservoir networks in an open loop. This gives $\delta_{k_{e}}$ of 10 LT. After this, we start autonomous closed-loop predictions with $\delta_{k_{e}}=10$. The value of $\delta_{k_{e}}$ represents the time difference from the start of closed-loop predictions to the extreme event. If the PH value is greater than $\delta_{k_{e}}$ we take PH equal to $\delta_{k_{e}}$. If the prediction diverges before the extreme event occurs, or the PH value is less than $\delta_{k_{e}}$, we decrease the $\delta_{k_{e}}$ value by a factor of $\tau_{e}$. We take $\tau_{e}$ = 0.5 LT and the new $\delta_{k_{e}}$ value equals to 9.5 LT. For the next step, we repeat the closed-loop predictions for the extreme event occurring at 9.5 LT, i.e., $\delta_{k_{e}}$ = 9.5 LT. We repeat the same method decreasing $\delta_{k_{e}}$ by $\tau_{e}$ until the PH is greater than $\delta_{k_{e}}$ (Fig. 17). That is, the extreme event is predicted correctly by the reservoir computers by the value of PH in the future.
@@
We calculate the PH for a set of reservoir sizes for both quantum and classical reservoir networks. For a classical reservoir, the number of neurons is varied from $N=256$ to $N=2048$. We tune the hyperparameter for each reservoir and calculate the corresponding PH value. We also perform the same analysis for the quantum reservoir with qubits $n=8$ to $n=11$, which, in principle, corresponds to the classical reservoirs of $N=256$ to $N=2048$. The performance of the quantum reservoir also improves with the reservoir size. In Fig. 18, the comparison of classical and quantum reservoirs for various reservoir sizes is shown. Both reservoir networks have comparable performances for reservoir sizes $N=\{256,512\}$, with the classical reservoir out-performing quantum reservoir by a small margin. However, the PH value of the classical reservoir converges with a median value of around 5.5 LT. This convergence of PH value with increasing reservoir sizes is due to performance saturation in CRC and it is consistent with the results in Ref. [18]. In contrast, the quantum reservoir outperforms the classical reservoir for increasing reservoir sizes and has a median PH value of 8.09LT with 11 qubits.
Figure 18: MFE system. Predictability Horizon (PH) for predictions with different reservoir sizes for the best performing CRC and QRC-C4 (RF-QRC) networks.
A second statistical measure to assess the extreme event prediction capabilities is the $F$-Score [57], which measures the model’s accuracy by combining metrics such as precision and recall. To calculate the $F$-score, we measure three different occurrences: (a) an extreme event occurs in the true data set and the model also predicts the extreme event (True Positive, TP) (b) there is no extreme event in the true data but the model predicts an extreme event (False Positive, FP) (c) an extreme event occurs in the true data set but the model does not predict an extreme event, (False Negative, FN) [18]. These three metrics are combined to calculate Precision ($p$), Recall ($r$), and $F$-score ($F$) from 1000 different starting points, which span 20 different time series
@@
Second, we fix the prediction time to 3 LT (PT = 3LT), which indicates the model is predicting extreme events between 3 - 4 LT. After fixing the PT value, we vary the reservoir sizes from $N=256$ to $N=2048$ and compare the performance metrics. The performances are similar for both quantum and classical reservoirs with reservoir sizes of $\{256,512\}$. Similar to the results shown for the predictability horizon previously, the quantum architecture has a higher $F$-score value for 10,11 qubits as compared to the classical reservoirs of the same sizes (Fig. 19). All the results are computed for 5 different ensembles of networks and for 1000 different starting points.
Figure 19: MFE system. $75^{th}$ percentile of Precision, Recall, F-score vs On the left side, Prediction Time (PT) for a constant reservoir size $(N=1024)$. On the right side, Reservoir Size for a constant Prediction Time $(PT=3LT)$.